A STATE SPACE MODEL FOR SOFTWARE RELIABILITY

YEN-CHANG CHANG AND LII-YUH LEU

Graduate Institute of Statistics, National Central University, Chung-Li, Taiwan, Republic of China

(Received July 1, 1996; revised October 13, 1997)

Abstract. In this paper, we propose a non-Gaussion state space model to apply in software reliability. This model assumes an exponential distribution for the failure time in every test-debugging stage, conditionally on the state parameter—the number of faults in the program. It is a generalized JM model which can be applied to the imperfect debugging situation as well as in evolving programs. By examining a set of data on evolving program failures, the effect of evolving program model is amply proved.

Key words and phrases: JM model, linear exponential loss function, Poisson distribution, exponential distribution.